OpenStudy (anonymous):

To win a prize, a person must select 3 numbers from 13 numbers. Find the probability of winning if a person buys one ticket. (Note: The numbers can be selected in any order.) Round to four decimal places, if necessary.

OpenStudy (anonymous):

i dont understand what you are asking me

OpenStudy (anonymous):

Ok, what would be the probability for picking 1 number of 13 correct? 1/13 right?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

So what would be the probability of, after picking the first number picking another number out of 13 right? \[\frac{1}{13} \times\frac{1}{13}\] right?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

So following the trend what would be the probability of picking 3 correct numbers out of 13?

OpenStudy (anonymous):

1/13

OpenStudy (anonymous):

So the probability for picking 1/13 numbers 3 times would then be...

OpenStudy (anonymous):

3/13

OpenStudy (anonymous):

that would be adding. Probability is generally multiplication. three coin flips in a row, dice it would be (1/13)(1/13)(1/13)

OpenStudy (anonymous):

do i multiply the denominator or keep it the same

OpenStudy (anonymous):

multiply

OpenStudy (anonymous):

so it would be 1/39

OpenStudy (anonymous):

yes in decimal form rounded to 4 digits. Think of it like this. If I took a dart board and broke it into fourths, the probability for hitting 1 of the fourths is 1/4 right? If I then cut each of the fourths into fourths and then chose only of the elements to hit the probability would be 1/16 since there are a total of 16 pieces. (1/4)(1/4) = (1/16)

OpenStudy (anonymous):

ok thank you so much i really do appreciate it mr. FREEWILLY922 LOL